Main INDEX
Monthly INDEX
PREV
NEXT
User name Y. Ilieva
Log entry time 23:58:03 on October 7,2006
Entry number 186563
keyword=swing shift summary
16:00 Swing shift begins. No beam.
16:15 Communication to iocha16 lost - ioc rebooted and the right NMR
probe rebooted several times without success.
Efforts to reboot the iocha16 and get the communication back continued till A. Camsonne found out there was a problem with
the display of hacsbc2.
18:10 hacsbc2 rebooted, but the problem with the display remains.
Hall A Main Menue and Hall A General Tools run on adaql8.
18:10 Q1 power has to be reset. Controlled access requested from MCC.
18:20 Hall A in controlled access. Q1 reset from the power supply rack.
18:31 Experiment ready again.
18:35 Beam permit requested from MCC.
19:27 Momentum reading of the Right Spectrometer dipole is out of spec. No control on the dipole's power supply. iocha16 and the right NMR power rebooted several times without any success.
20:01 The right dipole still not responding. Mark Stevens called.
He needs to come in to reset the power.
20:02 Controlled access requested from MCC.
20:17 Hall-A in controlled access, full survey begins.
20:40 Full survey completed.
20:42 M. Stevens in the Hall. Right dipole current ramped down.
22:10 Left dipole current ramping up.
At around 20:30 beam was delivered to other Halls.
around 22:20 Hall A back in Beam Permit.
23:33 Asked for 30 muA beam (dipole current close to set point)
in ordfer to check its quality.
Beam back, beam position out of spec.
23:56 Still adjusting the current on the right dipole to get it to
central momentum of 1.5607 GeV/c.